An 80N wooden block is placed on top of a wooden table. If a force of 28N is necessary to keep the block moving at a constant velocity. (a) What is the coefficient of sliding friction for the block & table. (b) If a weight of 40N is paced on top of the wooden block. What force will be required to keep the block and weight moving at a constant velocity across the table?
1
Expert's answer
2019-10-22T10:24:10-0400

(a)


"\mu=\frac{F}{W}=\frac{28}{80}=0.35"

(b)


"F=\mu(W+W')=0.35(80+40)=42\ N"